OpenStudy (anonymous):

Simplify: √-6(2+√-8)

OpenStudy (anonymous):

help please :(((

OpenStudy (whpalmer4):

\[\sqrt{-n} = \sqrt{-1*n} = i \sqrt{n}\]where\[i=\sqrt{-1}\text{ and } i^2=-1\]Using that, you can rewrite your expression as square roots of positive numbers and simplify it. Or, just use the distributive property first, and then apply the knowledge about \(i\). Give it a try and show me your work if you get stuck.

OpenStudy (anonymous):

So how would I rewrite it as square roots of positive numbers?

OpenStudy (anonymous):

If I distribut' would it look like √-12(√-56) ??

OpenStudy (anonymous):

GTG!!

OpenStudy (anonymous):

\[i \sqrt6(2+2 i \sqrt2)\] is a start

OpenStudy (whpalmer4):

Here's the distribution: \[\sqrt{-6}(2+\sqrt{-8}) = \sqrt{-6} * 2 + \sqrt{-6}*\sqrt{-8}\]Now use \[\sqrt{a}*\sqrt{b} = \sqrt{a*b}\]and\[\sqrt{-n} = i\sqrt{n}\]